OpenStudy (anonymous):

Tim paddles his kayak 12 km upstream against a 3km/h current and back again in 5 hours and 20 minutes. In that time, how far could he have paddled in still water?

OpenStudy (anonymous):

d=rt

OpenStudy (anonymous):

let the speed in still water be v say. time taken to go upstream t1 = 12/ v - 3 time taken to come downstream t2 = 12/ v + 3 t1 + t2 = 5 hours and 20 minutes --> 16/3 hours (12/ v - 3) + (12/ v + 3) = 16/3 solve this to find v. u will get v= 6 km/h (solve and recheck urself :) ) after finding v multiply the velocity with total time 16/3 hours to get distance covered in still water. that is 6*16/3= 32km this is a hard question but i think this should be right